If a random variable is normally distributed with mean 4.35 and standard deviation 0.59, what is the probability that this random variable will take on values more than 5.2?


1
Expert's answer
2021-05-02T13:03:26-0400

"\mu = 4.35 \\

\sigma = 0.59 \\

P(X>5.2) = 1 -P(X<5.2) \\

= 1 -P(Z< \frac{5.2-4.35}{0.59}) \\

= 1 -P(Z<1.44) \\

= 1 -0.9250 \\

= 0.0750"